OpenStudy (anonymous):

What is the equation of the line which includes points (0, 3) and (4, 5)? Please help!

OpenStudy (anonymous):

gradient = y2-y1/x2-x1

OpenStudy (anonymous):

5-3/4-0

OpenStudy (anonymous):

1/2 is the answer

OpenStudy (gorv):

y-y1= slope(x-x1)

OpenStudy (gorv):

slope you calculated as = 1/2 and put in any point in place of x1 and y1

OpenStudy (mathstudent55):

The slope-intercept equation of a line is: \(y = mx + b\)

OpenStudy (mathstudent55):

In this problem we are given two points, (0, 3) and (4, 5), so we can find the slope, m. \(m = \dfrac{y_2 - y_1}{x_1 - x_1} = \dfrac{5 - 3}{4 - 0} = \dfrac{2}{4} = \dfrac{1}{2}\)

OpenStudy (anonymous):

i already said that :/

OpenStudy (mathstudent55):

Now we can add the slope to the slope-intercept form of the equation. \(y = \dfrac{1}{2}x + b\) We need to find b.

OpenStudy (mathstudent55):

Now we use either one of our given points, and we plug in the x- and y-coordinates for x and y, respectively, in the equation to find b. Let's use point (0, 3): \(3 = \dfrac{1}{2}(0) + b\) \(b = 3\) Now that we know b = 3, we just replace b with 3 in the slope-intercept equation: \(y = \dfrac{1}{2}x + 3\)

OpenStudy (mathstudent55):

@Gaelle_1 You did calculate the slope correctly, but you stopped there, and said that was the answer. The question is asking for the equation of the line, not just the slope.

OpenStudy (anonymous):

Actually, the answer was indeed only y=1/2x plus 3 I forgot about this and got it lol

OpenStudy (anonymous):

@mathstudent55 yeah sorry :P